Database last updated on Sat, 31 Jan 2026 19:10:53 +0000 by localhost

Instrument constant changes for the week before 2024042 (2 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 17 rows)

Measurements and tests summary.


_jday_loc_serial_skcdtedhgrsslzs
042TORONTO#017cal,Next schedule will be cal (now cal ZS= 85 ZS= 85.21499),The next day will run schedule cal,Waiting until end of local day.,Waiting until ZA =32934103
(0 sec)

Number of days since last dome/window cleaning (0 sec, 2 rows)

serialdays
017more than a year

Missing files (0 sec, 1 rows)

Seriallast datadays ago
01720240430-1

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
0172423010210

AP results for MB Brewer. (0 sec, 1 rows)

SerialJdayt1 offt1 ont2 offt2 ont3 offt3 onHV offHV on+15V off+15V on+5V off+5V on-15V off-15V on+24V off+24V on+5V sps off+5V sps on-8V sps off-8V sps onSL A offSL A onSL V offSL V onW
01720240425.96.07.17.16.76.61318.81304.215.114.85.15.1-15.0-15.024.224.34.94.9-8.0-8.09.11.60.010.3016.2

Suspicious HG test results (correlation<0.98 or counts<5000). (0 sec, 1 rows)

SerialJdaycountPeakmin.corr.max.corr.30-day history
#0170421500680.97870.9787